OpenStudy (anonymous):

Use the Pythagorean Theorem to find the length of the missing side of a right triangle whose legs equal √2 and 2√15. a right triangle with legs as 2 square root of 15 and square root of 2. A. 64 B. 4 + 2√15 C. 8 D. None of the above

OpenStudy (whpalmer4):

Well, by Pythagorean Theorem, the square of the hypotenuse is the sum of the squares of the two legs: \[a^2+b^2=c^2\] for legs a,b and hypotenuse c You have \(a=\sqrt{2}\) and \(b = 2\sqrt{15}\) (or vice versa), what do you get when you square them and add them together? Take the square root of the result to find the length of the hypotenuse.

OpenStudy (whpalmer4):

I would call this a bit of a trick question, so do your work carefully!